OpenStudy (anonymous):

84.41 is 23% of what number?

OpenStudy (gw2011):

0.23x=84.41 x=84.41/0.23 x=367

OpenStudy (anonymous):

Let the unknown number be x. 0.23*x=84.41 84.41/0.23=x x=367